link co just paid a dividend of $1.00 per share. Analysts expect its dividend to grow at 20% per year for the next two years and then 3% per year thereafter. If the required rate of return in the stock is 7%, calculate the current value of the stock.

$32.49

$34.77

$35.82

$36.00

$37.20
Answer #1

D1=(1*1.2)=1.2

D2=(1.2*1.2)=1.44

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(1.44*1.03)/(0.07-0.03)

=37.08

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.2/1.07+1.44/1.07^2+37.08/1.07^2

=$34.77(Approx).